# Weakly-Supervised Conditional Embedding for Referred Visual Search

### v2

@@ 5.1 Implementation details
All our models take as input images of size $224\times 224$, and output an embedding vector of 512 dimensions. We use CLIP weights as initialization, and then train our models for 30 epochs with AdamW (Loshchilov & Hutter, 2019) and a maximum learning rate of $10^{-5}$ determined by a learning rate range test (Smith, 2017). To avoid distorting pretrained features (Kumar et al., 2022), we start by only training the final projection and new input embeddings (conditioning and positional) for a single epoch, with a linear warm-up schedule. We then train all parameters for the rest of the epochs with a cosine schedule.
@@ 5.2 Results
We leveraged the recent Grounding DINO (Liu et al., 2023) and Segment Anything (Kirillov et al., 2023) to create a baseline approach based on object detection and segmentation. In this setting, we feed the model the query image and conditioning information, which can be either the name of the category or a caption. Subsequently, we use the output crops or masks to train a ViT following the aforementioned procedure. Please refer to Tab. 1 for the results.
@@
We also compare with ASEN (Dong et al., 2021) trained on our dataset using the authors’ released code. This conditional architecture uses a global and a local branch with conditional spatial attention modules, respectively based on ResNet50 and ResNet34 backbones, with explicit ROI cropping. However in our experiments the performances decrease with the addition of the local branch in the second training stage, even after tuning the hyperparameters. We report results for the global branch.
@@
To further validate our approach, we replaced the categorical conditioning with referring expressions, using our generated BLIP2 captions embedded by a Sentence T5-XL model (Ni et al., 2022). We chose this model because it embeds the sentences in a 768-dimensional vector, allowing us to simply replace the categorical token. We pre-computed the caption embeddings, and randomly used one of them instead of the product category at training time. At test time, we used the first caption.
@@
We compare these models with CIR methods: CoSMo (Lee et al., 2021) and CLIP4CIR (Baldrati et al., 2023). Both use a compositor network to fuse features extracted from the image and accompanying text. CoSMo reaches performances similar to an unconditional ViT-B/32, while CLIP4CIR performs similarly to our textual CondViT-B/32. We hypothesize that for our conditional feature extraction task, early conditioning is more effective than modifying embeddings through a compositor at the network’s end. Our CondViT-B/16 model significantly outperforms all other models and achieves results comparable to our caption-based approach using Grounding DINO-B (see Fig. 4, right). As the RVS task differs from CIR, despite both utilizing identical inputs, this was anticipated. Importantly, CondViT-B/16 accomplishes this without the need for explicit detection steps or dataset-specific preprocessing. Notably, we observe that our models achieve a categorical accuracy of $98\%$ against 2M distractors, surpassing the accuracy of the best corresponding detection-based model, which stands at $94.3\%$.
@@
Figure 5: Qualitative results for our categorical (first 2 rows) and textual (last 2 rows) CondViT-B/16. We use free-form textual queries instead of BLIP2 captions to illustrate realistic user behavior, and retrieve from the whole test gallery. See Fig. 8 and 9 in the Appendix for more qualitative results.
@@ 6 Conclusion & Limitations
7 Ethics Statement
Harmful and Private Content.
Dataset Biases.
Retrieval Systems.
The broader impact of this research is similar to other works on instance retrieval, with the significant advantage of producing embeddings that conceal information about personal identity and other undesired elements thanks to the conditioning.
8 Reproducibility Statement
The code used to produce the results will be released. Additionally, our training setup is described in Section 5.1.
The dataset will be publicly released. The method we used to create it is thoroughly described in Appendix A.2.
@@ A.1 Samples
A.2 Construction
@@
We used a second model to automatically assign categories to the simple images. LAION captions are noisy, so instead of using them we used BLIP2 FlanT5-XL (Li et al., 2023) to answer the question "In one word, what is this object?". We gathered all the nouns from the answers, using POS tagging when the generated answer was longer, and grouped them in 11 categories (10 for clothing, 1 for non-clothing). We automatically created an initial pool of labeled data, which we manually filtered, before applying the same active learning process as above. We then annotated all the simple images with this model. Please refer to Appendix A.3 for the list of categories and their composition.
@@
Finally, we used Locally Sensitive Hashing (LSH) with perceptual hash, and OpenCLIP B16 embeddings to remove duplicates. We created FAISS indexes based respectively on hamming distance and cosine similarity, automatically removing samples with extremely high similarity. We manually inspected samples near the threshold. We used this process on complex images from the training set to remove products duplicates, on train and test sets to reduce evaluation bias, and on gallery images and distractors for both the validation and test sets.
A.3 Composition
@@ Appendix B Model
B.1 Retrieval Examples
In this section, we show additional results for our categorical CondViT-B/16 and its textual variant trained with BLIP2 (Li et al., 2023) captions. We use test query images and the full test gallery with 2M distractors for the retrieval. Each query in the test set is exclusively associated with a single item. However, it should be noted that the we do not necessarily query for this item, so the queried product might not be in the gallery. Nevertheless, owing to the presence of 2M distractors, most queries can retrieve multiple viable candidates.
@@ B.1 Retrieval Examples
Figure 9: Retrieved items for queries in LRVS-F test set with our textual CondViT-B/16. 9(a) shows results for normal, concise use. 9(b) shows results with more verbose queries. 9(c) shows queries influencing the type of results. 9(d) show results for out-of-frame items.
B.2 Attention Maps
ViT-B/16 ASEN Cat. CondViT-B/16
Upper Body Lower Body Bags Upper Body Lower Body Bags
Upper Body Lower Body Feet Upper Body Lower Body Feet
Upper Body Lower Body Feet Upper Body Lower Body Feet
Upper Body Lower Body Feet Whole Body Lower Body Bags
Upper Body Lower Body Bags Upper Body Lower Body Bags
Figure 11 shows the patch features extracted by our models with principal component analysis (PCA) computed on all image tokens in the last layer of our CondViT-B/16 model across the test queries. Similarly to Oquab et al. (2023), we find that applying a threshold on the first component enables effective separation of the background from the foreground. Intriguingly, we observe that employing a higher threshold not only accomplishes the aforementioned separation but also yields cleaner visualizations by isolating the conditionally selected object. We also observe instances where the network encounters difficulties in detecting the referenced object, resulting in a notable absence of tokens surpassing the established threshold.
B.3 Ablation Studies
Insertion Depth.
Asymetric Conditioning.
Distractors $\rightarrow$ +0 +10K +100K +1M +2M
Model %R@1 %Cat@1 %R@1 %Cat@1 %R@1 %Cat@1 %R@1 %Cat@1 %R@1 %Cat@1
B.4 Textual Conditioning — Failure Cases
"A red handbag" "A green stripped t-shirt"
"Pants" "A scarf"
Figure 13: Retrieved items showing failure cases of our textual CondViT-B/16. 13(a) shows that the network disregards color clues. 13(b) shows that the network defaults to the salient item when the query is too hard or not represented.
@@ C.1 Motivations
    This dataset has been created to provide public training data and a benchmark for the Referred Visual Search (RVS) task, for research purposes. The task is new, and thereby no other dataset existed to tackle it.
@@
    [Anonymized].
@@ C.2 Composition
    Instances of this dataset are URLs from online catalogs of fashion retailers, and a such they depict either fashion products or models wearing them.
@@
    We added mutiple synthetic labels to the images. First, a type, Complex when the image depicts a scene, with a model, Simple when it is an isolated product. There also exist a Partial Complex category, for scene images that are zoomed-in and do not contain the entire product. Second, a product identifier, allowing to group images depicting the same product. Each simple target image is further described by a category, following the taxonomy described in this paper, and 2 BLIP2-FlanT5XL captions.
@@
    Yes, complex images often depict multiple objects, but are linked with only one product in this dataset. They are registered in fashion catalogs with the intent to showcase a specific product, and as such we were not able to extract more information.
@@
    Yes, we provide a synthetic product identifier for each image, allowing to group simple and complex images depicting the same product.
@@
    Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
@@
    It might be possible to identify models using face recognition, but it would require external data.
@@ C.3 Collection Process
    We manually curated domains and manually designed regular expressions to extract product identifiers from the URLs. The additional labels and captions are synthetic. We validated the quality of the labels by mesuring accuracy on random samples, and the captions with a CLIP similarity. Most of the process was done on a single CPU node, with the exception of the deep learning models which were run on two GPUs.
@@
    The authors are the only persons involved in this data collection process.
@@
    This dataset and LAION 5B have been filtered using CLIP-based models. They inherit various biases contained in their original training set. Furthermore, the selected domains in this work only represent European and American fashion brands, and do not provide
@@ C.4 Preprocessing / Cleaning / Labeling
    Please refer to Appendix. A.2 for additional details.
@@
    The raw data is LAION-5B. It is available through LAION’s HuggingFace pages.
@@ C.6 Distribution
    We release our data under the MIT license.
@@ C.7 Maintenance
    [Anonymized]
@@
    There is no erratum as this is the initial release. If need be, we will update the dataset page on [Anonymized].
@@
    We will not update the dataset, as it contains a benchmark and we want the results to stay comparable across time.
@@
    We do not plan on supporting extensions to this dataset as it is intended to be a benchmark and results must stay comparable across time. However we do encourage the creation of new similar datasets across new verticals, to extend the field of Referred Visual Search.

### v3

@@ 5.1 Implementation details
All our models take as input images of size $224\times 224$, and output an embedding vector of 512 dimensions. We use CLIP weights as initialization, and then train our models for 30 epochs with AdamW [33] and a maximum learning rate of $10^{-5}$ determined by a learning rate range test [43]. To avoid distorting pretrained features [26], we start by only training the final projection and new input embeddings (conditioning and positional) for a single epoch, with a linear warm-up schedule. We then train all parameters for the rest of the epochs with a cosine schedule.
@@ 5.2 Results
We leveraged the recent Grounding DINO [30] and Segment Anything [24] to create a baseline approach based on object detection and segmentation. In this setting, we feed the model the query image and conditioning information, which can be either the name of the category or a caption. Subsequently, we use the output crops or masks to train a ViT following the aforementioned procedure. Please refer to Tab. 1 for the results.
@@
We also compare with ASEN [8] trained on our dataset using the authors’ released code. This conditional architecture uses a global and a local branch with conditional spatial attention modules, respectively based on ResNet50 and ResNet34 backbones, with explicit ROI cropping. However in our experiments the performances decrease with the addition of the local branch in the second training stage, even after tuning the hyperparameters. We report results for the global branch.
@@
To further validate our approach, we replaced the categorical conditioning with referring expressions, using our generated BLIP2 captions embedded by a Sentence T5-XL model [37]. We chose this model because it embeds the sentences in a 768-dimensional vector, allowing us to simply replace the categorical token. We pre-computed the caption embeddings, and randomly used one of them instead of the product category at training time. At test time, we used the first caption.
@@
We compare these models with CIR methods: CoSMo [28] and CLIP4CIR [2]. Both use a compositor network to fuse features extracted from the image and accompanying text. CoSMo reaches performances similar to an unconditional ViT-B/32, while CLIP4CIR performs similarly to our textual CondViT-B/32. We hypothesize that for our conditional feature extraction task, early conditioning is more effective than modifying embeddings through a compositor at the network’s end. Our CondViT-B/16 model significantly outperforms all other models and achieves results comparable to our caption-based approach using Grounding DINO-B (see Fig. 5, right). As the RVS task differs from CIR, despite both utilizing identical inputs, this was anticipated. Importantly, CondViT-B/16 accomplishes this without the need for explicit detection steps or dataset-specific preprocessing. Notably, we observe that our models achieve a categorical accuracy of $98\%$ against 2M distractors, surpassing the accuracy of the best corresponding detection-based model, which stands at $94.3\%$.
@@
Figure 6: Qualitative results for our categorical (first 2 rows) and textual (last 2 rows) CondViT-B/16. We use free-form textual queries instead of BLIP2 captions to illustrate realistic user behavior, and retrieve from the whole test gallery. See Fig. 13 and 14 in the Appendix for more qualitative results.
@@ 6 Conclusion & Limitations
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes]
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix A.5.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Appendix A.5.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Appendix A.2.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 5.1.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Table 1 and 2.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 5.1.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We provide the URLs to the assets in Appendix A.2.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ A.1 Samples
A.2 Usage
The dataset is hosted on the HuggingfaceHub, using the widely used parquet format, released under the CC-BY-NC-4.0 license, for research use only. The images can easily be downloaded using tools like img2dataset.
A.3 Construction
@@
We used a second model to automatically assign categories to the simple images. LAION captions are noisy, so instead of using them we used BLIP2 FlanT5-XL [29] to answer the question "In one word, what is this object?". We gathered all the nouns from the answers, using POS tagging when the generated answer was longer, and grouped them in 11 categories (10 for clothing, 1 for non-clothing). We automatically created an initial pool of labeled data, which we manually filtered, before applying the same active learning process as above. We then annotated all the simple images with this model. Please refer to Appendix A.4 for the list of categories and their composition.
@@
Finally, we used Locality Sensitive Hashing (LSH) with perceptual hash, and OpenCLIP B16 embeddings to remove duplicates. We created FAISS indexes based respectively on hamming distance and cosine similarity, automatically removing samples with extremely high similarity. We manually inspected samples near the threshold. We used this process on complex images from the training set to remove products duplicates, on train and test sets to reduce evaluation bias, and on gallery images and distractors for both the validation and test sets.
A.4 Composition
@@
A.5 Ethics Statement
Harmful and Private Content.
Dataset Biases.
@@ Appendix B Model
B.1 Ablation Studies
Insertion Depth.
Asymetric Conditioning.
Distractors $\rightarrow$ +0 +10K +100K +1M +2M
Model %R@1 %Cat@1 %R@1 %Cat@1 %R@1 %Cat@1 %R@1 %Cat@1 %R@1 %Cat@1
B.2 Attention Maps
ViT-B/16 ASEN Cat. CondViT-B/16
Upper Body Lower Body Bags Upper Body Lower Body Bags
Upper Body Lower Body Feet Upper Body Lower Body Feet
Upper Body Lower Body Feet Upper Body Lower Body Feet
Upper Body Lower Body Feet Whole Body Lower Body Bags
Upper Body Lower Body Bags Upper Body Lower Body Bags
Figure 11 shows the patch features extracted by our models with principal component analysis (PCA) computed on all image tokens in the last layer of our CondViT-B/16 model across the test queries. Similarly to Oquab et al. [39], we find that applying a threshold on the first component enables effective separation of the background from the foreground. Intriguingly, we observe that employing a higher threshold not only accomplishes the aforementioned separation but also yields cleaner visualizations by isolating the conditionally selected object. We also observe instances where the network encounters difficulties in detecting the referenced object, resulting in a notable absence of tokens surpassing the established threshold.
B.3 Textual Conditioning — Failure Cases
"A red handbag" "A green stripped t-shirt"
"Pants" "A scarf"
Figure 12: Retrieved items showing failure cases of our textual CondViT-B/16. 12(a) shows that the network disregards color clues. 12(b) shows that the network defaults to the salient item when the query is too hard or not represented.
B.4 Retrieval Examples
In this section, we show additional results for our categorical CondViT-B/16 and its textual variant trained with BLIP2 [29] captions. We use test query images and the full test gallery with 2M distractors for the retrieval. Each query in the test set is exclusively associated with a single item. However, it should be noted that the we do not necessarily query for this item, so the queried product might not be in the gallery. Nevertheless, owing to the presence of 2M distractors, most queries can retrieve multiple viable candidates.
@@ B.4 Retrieval Examples
Figure 14: Retrieved items for queries in LRVS-F test set with our textual CondViT-B/16. 14(a) shows results for normal, concise use. 14(b) shows results with more verbose queries. 14(c) shows queries influencing the type of results. 14(d) show results for out-of-frame items.
@@ C.1 Motivations
    This dataset has been created to provide public training data and a benchmark for the Referred Visual Search (RVS) task, for research purposes. The task is new in academic research, and thereby no other dataset existed to tackle it.
@@
    The dataset was created by Simon Lepage, Jérémie Mary and David Picard, on behalf of the CRITEO AI Lab and ENPC.
@@ C.2 Composition
    Instances of this dataset are URLs from online catalogs of fashion retailers. The associated images depict fashion products isolated and in context. Some products are associated with a synthetic product identifier.
@@
    We added multiple synthetic labels to the images. First, a type, Complex when the image depicts a scene, with a model, Simple when it is an isolated product. There also exist a Partial Complex category, for scene images that are zoomed-in and do not contain the entire product. Second, a product identifier, allowing to group images depicting the same product. Each simple target image is further described by a category, following the taxonomy described in this paper, and 2 BLIP2-FlanT5XL captions.
@@
    Yes, complex images often depict multiple objects, but are linked with only one product in this dataset. They are registered in online fashion catalogs with the intent to showcase a specific product, and as such we were not able to extract more information.
@@
    Yes, we provide a synthetic product identifier for each image (exluding distractors), allowing to group simple and complex images depicting the same product.
@@
    Does the dataset identify any subpopulation (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
@@
    It might be possible to identify models using facial recognition, but it would require external data.
@@ C.3 Collection Process
    We manually curated domains and manually designed regular expressions to extract product identifiers from the URLs. The additional labels and captions are synthetic. We validated the quality of the labels by measuring accuracy on random samples, and the captions with a CLIP similarity. Most of the process was done on a single CPU node, with the exception of the deep learning models which were run on two GPUs.
@@
    The authors were the only persons involved in this data collection process.
@@
    This dataset and LAION 5B have been filtered using CLIP-based models. They inherit various biases contained in their original training set. Furthermore, the selected domains in this work only represent European and American fashion brands, and do not provide a comprehensive view of worldwide fashion.
@@ C.4 Preprocessing / Cleaning / Labeling
    Please refer to Section. 3.1 and Appendix A.3 for additional details.
@@
    The "raw" data is LAION-5B.
@@ C.6 Distribution
    We release our data under the CC-BY-NC-4.0 license.
@@ C.7 Maintenance
    The owner of the dataset can be contacted through the dataset’s HuggingFace space.
@@
    There is no erratum as this is the initial release. If need be, we will update the dataset repository.
@@
    We do not plan to update the dataset, as it contains a benchmark and we want the results to stay comparable across time.
@@
    We do not plan on supporting extensions to this dataset as it is intended to be a benchmark and results must stay comparable across time. However we do encourage the creation of similar datasets across new verticals, to extend the field of Referred Visual Search.